OpenStudy (anonymous):

Solve for x: 3x - 5 = 2x + 6.

OpenStudy (anonymous):

3x-5=2x+6 subtract 2x from both sides x-5=6 add 5 on both sides x=11

OpenStudy (anonymous):

Solve for x: -3(x + 3) = -3(x + 1) - 5. 6 -6 All real numbers No solution

OpenStudy (anonymous):

no solution